# Asymptotic constraints for 1D planar grey photon diffusion from linear transport with special-relativistic effects

### v1

@@ 1 Introduction
Radiation (photon, neutrino) diffusion through optically thick matter moving at (near-)relativistic speed or in the presence of gravity effects is a ubiquitous phenomenon in the study of astrophysical plasmas. It is well known to be a limiting behavior of the more general phenomenon of radiative transfer. For instance, photon propagation through expanding supernova ejecta can be characterized with a diffusion solution (see, for instance, [arnett1982, pinto2000]). Numerical implementation of the diffusion equation can be used to accelerate radiative transfer in optically thick regions of space and frequency. One example is Monte Carlo (MC) accelerated with random walk (RW) or discrete diffusion Monte Carlo (DDMC) [fleck1984, gentile2001, densmore2007, densmore2012, cleveland2014, keady2017, smith2020]. These RW and DDMC methods have been developed for applications to: neutrinos in core collapse supernova engines [abdikamalov2012], cosmic rays [harding2016, fitzaxen2021], radiation pressure feedback in stellar cluster dynamics [tsang2018], supernova and kilonova spectral synthesis [wollaeger2013, wollaeger2014, wagle2023], and Lyman $\alpha$ photon transfer during galaxy evolution [smith2018].
While diffusion theory has found practical use in describing photon and neutrino fields in astrophysical plasmas, the infinite signal speed of the diffusion equation is incompatible with special relativity (see, for instance, [dunkel2007, gzyl2025]). This can readily be seen for static material backgrounds, where the Green’s function conforms to a Gaussian distribution, which has non-zero values everywhere. As a result, the applicability of diffusion is limited to high-optical depth regions, unless the equation is modified to restrict the signal speed (for instance $P_{1}$, $M_{1}$ [levermore1984] and flux-limit diffusion [levermore1981]). We may also see the effect of Lorentz transforming the space-time derivatives in the diffusion equation. Suppose that a frame of reference, denoted by subscript 0, possesses particles that adhere to a 1D diffusion equation,
@@
where $t_{0}$ is time, $x_{0}$ is the spatial coordinate, $D$ is a constant diffusion coefficient, and $\phi_{0}$ is the particle density. Applying the standard Lorentz transformations for the partial derivatives (see, for instance, [lowrie2023]) gives
@@
Thus in the new frame, distinguished here by the absence of a subscript, a second derivative in time is manifest. Similar to the infinitely extended signal in space, this second derivative in time must couple all time from the past to the future and vice versa. Consequently, this equation is not immediately amenable to numerical schemes that can only advance forward in time in the frame of $(x,t)$.
@@
Stochastic analyses originated by [hakim1965] and [dudley1966] (and further developed in subsequent works) demonstrate that relativity breaks Markovian processes in space, where one would expect that the set of spatial positions for the current particle state completely determine the next state [dudley1966, dunkel2007, giona2017]. In developing the relativistic theory, one might first consider bounding the particle speed in a Poisson-Kac process [kac1974], which corresponds to the Telegrapher ($P_{1}$) equation and is non-Markovian [kac1974, dunkel2007]. Spatially non-Markovian stochastic processes have been formulated in special [dunkel2007, serva2020] and general [herrmann2010] relativistic settings. Properly accounting for relativity in a collective sense, where a frame is moving relative to the frame in which diffusion is a good description, is a somewhat distinct problem. To this end, [giona2017] examine the particle motion as a Poisson-Kac process along discrete directions, where the process exists in its usual form in one inertial (comoving) frame, finding in 1D that the effective diffusion coefficient in another inertial frame is $1/\gamma^{3}$ times the comoving coefficient; [giona2017] observe that this follows the intuition that $D\sim\delta x^{2}/\delta t\sim(\delta x_{0}/\gamma)^{2}/(\gamma\delta t_{0})$, from Lorentz contraction of space and time dilation.
@@
For asymptotic analysis, the principal assumption is typically that the inverse of the macroscopic opacity is large compared to all length scales of interest [thomas1930]. Unlike the works above that focus on the statistical mechanics of the particles, this approach usually: starts from the full radiative transfer equation, imposes the assumed parameter scalings, and expands the radiation intensity in powers of a small-scale parameter (sometimes labeled $\varepsilon$). Asymptotic analysis has been applied to deriving a semi-relativistic diffusion equation [morel2006, lowrie2014], where $\beta$ is assumed to be asymptotically small (equivalently: the hydrodynamic flow time scale is assumed much longer than the light crossing time at the length scale of interest). In the fully relativistic context, including general relativity, asymptotic analysis has been examined early on by [thomas1930] and subsequently generalized by [anderson1972] (see also [achterberg2018]). Building off of these works, [shibata2011] demonstrate the truncated Thorne moment formalism (see [thorne1981]) converges to a diffusion equation with modifications from fluid acceleration and shear in the optically thick limit. The analyses of [anderson1972] and [shibata2011] consider both asymptotically large absorption and scattering opacity. This is consistent with the static-material scaling that furnishes equilibrium diffusion from non-linear thermal radiative transfer, but departs in not assuming that the light crossing distance on the time scale of interest is large compared to the spatial domain of interest [larsen1983].
@@
In the present study, we instead consider an asymptotic analysis that restricts to a special relativistic, 1D-planar, linear transport context, and we focus on the correspondence of the lab frame (some inertial frame) to the non-inertial frame comoving with fluid parcels of the background matter. In a departure from the work of [anderson1972] and [shibata2011] we attempt to use the standard parameter scaling for linear transport (see, for instance, [habetler1975, larsen1992]) in the comoving frame. Thus the scattering opacity is assumed to be asymptotically large while the absorption opacity is assumed to be asymptotically small, and there is not an equilibrium solution (for instance, the Planck or Fermi-Dirac distributions in the comoving frame). In doing so, our objective is to: (i) highlight apparent non-physical pathologies that occur from applying this standard scaling when velocity gradients are present, (ii) demonstrate that a particular scaling furnishes the Poisson-Kac-derived $\gamma$-dressed result of [giona2017], but for continuous directions and including the standard grey contribution from adiabatic expansion and Doppler shift (see, for instance [castor2004]), (iii) compare solutions of the diffusion equation to a relativistic linear transport solution and better-known semi-relativistic diffusion equation [castor2004].
@@
This paper is organized as follows. In Section 2, we present all considerations for deriving the fully relativistic 1D planar grey diffusion equation, starting from a linear transport equation with isotropic opacity and elastic isotropic scattering in the comoving frame. This includes considering the effect of frequency Doppler shift on integration over frequency for the lab-frame grey equation, in Section 2.1. We present properties of a simple class of functions, which are weighted powers of Lorentz transform (Doppler shift) factors of frequency, and their associated integrals in Section 2.2. These functions and their properties facilitate both the asymptotic derivation and the lab-frame harmonic expansion (Appendix AB). In Section 2.4, we present the development of the lab-frame asymptotic scaling that furnishes a closed comoving equation with respect to lab-frame space-time coordinates. This section is subdivided into sections that showcase the pathologies encountered from simpler asymptotic scaling attempts. In Section 3, we present a discretization of the fully relativistic diffusion equation, identifying terms with corresponding non- or semi-relativistic DDMC analogs. We then compare this discretization to lab-frame MC transport to provide a preliminary exploration of where the fully relativistic diffusion equation is applicable in Section 4. In the appendix, we provide: (A) details of the integrated Doppler function recursion, (B) the aforementioned harmonic ($P_{1}$) expansion, and finally (C) a brief non-relativistic attempt of the apparent opacity-only scaling of [thomas1930, anderson1972, shibata2011].
@@ 2.4 Asymptotic scaling
where $A[\cdot]$ is the original time-derivative operator (or the remainder of the transport equation), $B(\nu)$ is an as-yet unknown coefficient and $A_{\varepsilon}[\cdot]$ is a $\varepsilon$-scaled form of this operator. This scaling assumes that the residual excluded from the $\varepsilon$ coefficient scales as a leading-order contribution. The full set of scaling relations we use is
@@ 2.4.2 Scaling the comoving time derivative with the anisotropic term
We see that we have obtained a drift-diffusion equation for the leading-order scalar intensity using our lab-frame asymptotic scaling with $B(\mu)$ from Eq. (36). Furthermore, we note that we have not assumed a constraint on the value of $\beta$ itself. The diffusion stencil has three factors of $1/\gamma$, consistent with the finding of [giona2017] for bi-directional 1D planar random walk. However, we have an additional $\beta$-dependent coefficient of $\phi_{0}^{(0)}$, which accounts for the energetic effect of Doppler shift and adiabatic expansion. To more clearly see this, we may neglect the time-dependence of $\beta$, and neglect terms of O($\beta^{2}$) and O($\beta\partial\beta/\partial x$), to obtain
@@
which is the standard semi-relativistic form (see, for instance, [castor2004], Chapter 6). If $\beta$ is constant and approaches 1, assuming $q$ and $\phi_{0}^{(0)}$ remain bounded,
@@
which is the limiting behavior observed by [giona2017].
@@ 2.4.3 Scaling the Lagrangian time derivative of $\beta$
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{Q_{0}}{\gamma}\;\;.$ (52)
where we have replaced the integral of the comoving external source with $Q_{0}$. This equation now has the same coefficient for adiabatic and expansion effects as the standard semi-relativistic form, so retains mirrored behavior in scalar intensity when $(x,\beta)\rightarrow(-x,-\beta)$. In subsequent sections we examine Eq. (52), which corresponds to the following list of lab-frame scaling relations, now including Eq. (51),
@@ 3.1 Spatial stencil
Here we describe the discretization of Eq. (48) we use to explore properties of the solutions in Section 4. We solve the resulting discretized system deterministically, but we note that there exists a straightforward Discrete Diffusion Monte Carlo (DDMC) [gentile2001, densmore2007, abdikamalov2012] interpretation of the discretized terms, and write them following the notation of [densmore2007]. In 1D, a grey deterministic solution is straightforward to implement, and we do not seek to hybridize lab-frame MC transport with DDMC in the scope of this work.
@@
$\eta_{i}=\frac{4}{3}\left.\frac{\partial\beta}{\partial x}\right|_{i}$ (54)
@@
can be viewed as advection “leakage opacities” [densmore2007] (see [abdikamalov2012] for an operator-split continuous treatment of advection).
The diffusion operator terms can be evaluated in the usual way in the domain interior [densmore2007],
@@
where $\pm$-superscripts again indicate evaluation just to the lower or upper side of the cell edge $i+1/2$. The formulation permits several options for discretizing the velocity field, for instance: if $\gamma$ is continuous then $\gamma_{i+1/2}^{-}=\gamma_{i+1/2}^{+}=\gamma_{i+1/2}$; if $\gamma$ is piecewise-constant, then one may set $\gamma_{i+1/2}^{-}=\gamma_{i}$ and $\gamma_{i+1/2}^{+}=\gamma_{i+1}$. Consequently, the diffusion leakage opacities, corresponding to the standard non-relativistic form [densmore2007], are
@@
Finally, to treat the boundary, we assume the asymptotic boundary layer condition used by [densmore2007] holds in the comoving frame,
@@
where $W(\cdot)$ is the angular weighting function for intensity incident just below $x_{i-1/2}$, $\psi_{0}^{-}$, and $\lambda$ is the extrapolation distance [habetler1975]. Equation (63) is consistent with applying scaling Eqs. (47) to the transport equation, keeping terms up to O($\varepsilon$), transforming intensity and opacity to the comoving frame, and then applying the boundary layer analysis of [habetler1975] (the factor of $\gamma$ from the transformation of $\sigma_{t}$ to $\sigma_{t,0}$ can be included in the transformation of the spatial coordinate to the stretched optical depth parameter). Equation (63) can also be obtained from the the static boundary condition specified in comoving space-time, assuming the comoving time derivative is 0. Following the discretization procedure of [densmore2007], the leakage opacity and boundary transmission probability are
@@ 3.2 Time stencil and stability
As the right panel of Fig. 1 suggests, increasing the time-centering parameter makes the time-evolution of the error more stable; for the parameters considered the stability profile of the drift-diffusion equation is comparable to that of the static diffusion equation. Furthermore, $\alpha=0.5$ should furnish 2nd-order accuracy in time. In Section 4, we restrict to steady state for comparing accuracy of the diffusion equation to lab-frame Monte Carlo transport; thus we eliminate the issue of the time-stability considered in the present section.
@@ 4 Numerical results
where $d_{p}\in i$ implies the sum is over the set of tracks of particle $p$ inside cell $i$. We note that expanding the quadratic and distributing the sum furnishes path length estimators for lab-frame energy density, flux and pressure in the context of the standard energy-momentum tensor transformation for comoving energy density (see, for instance, [castor2004], Chapter 6).
For the following calculations, we fix the spatial domain $x\in[0,1]$ with 128 spatial cells. We also fix the comoving opacities as $\sigma_{t,0}=128$ and $\sigma_{a,0}=1/2$. All problems have vacuum boundary conditions on both sides of the spatial domain. Units are omitted as the linear transport and diffusion equations are spatially scale-free, so $\sigma_{t,0}=128$ and $\partial\beta/\partial x=0.03$ are 128 mean-free paths and a change in $\beta$ by 0.03 over some unit of length (e.g. the radius of a star). For the MC transport, we employ 512 particles per cell for the uniform volume source and 65,536 particles for the point source.

### v2

@@ 1 Introduction
Radiation (photon, neutrino) diffusion through optically thick matter moving at (near-)relativistic speed or in the presence of gravity effects is a ubiquitous phenomenon in the study of astrophysical plasmas. It is well known to be a limiting behavior of the more general phenomenon of radiative transfer. For instance, photon propagation through expanding supernova ejecta can be characterized with a diffusion solution (see, for instance, [4, 29]). Numerical implementation of the diffusion equation can be used to accelerate radiative transfer in optically thick regions of space and frequency. One example is Monte Carlo (MC) accelerated with random walk (RW) or discrete diffusion Monte Carlo (DDMC) [12, 13, 8, 7, 6, 21, 32]. These RW and DDMC methods have been developed for applications to: neutrinos in core collapse supernova engines [1], cosmic rays [18, 11], radiation pressure feedback in stellar cluster dynamics [36], supernova and kilonova spectral synthesis [38, 39, 37], and Lyman $\alpha$ photon transfer during galaxy evolution [33].
While diffusion theory has found practical use in describing photon and neutrino fields in astrophysical plasmas, the infinite signal speed of the diffusion equation is incompatible with special relativity (see, for instance, [10, 15]). This can readily be seen for static material backgrounds, where the Green’s function conforms to a Gaussian distribution, which has non-zero values everywhere. As a result, the applicability of diffusion is limited to high-optical depth regions, unless the equation is modified to restrict the signal speed (for instance $P_{1}$, $M_{1}$ [24] and flux-limit diffusion [25]).
It is straightforward to show that the diffusion equation is not Lorentz-invariant. For instance, consider the 1D planar diffusion equation,
@@
where $t_{0}$ is time, $x_{0}$ is the spatial coordinate, $D$ is a constant diffusion coefficient, and $\phi_{0}$ is the particle density. Applying the standard Lorentz transformations for the partial derivatives (see, for instance, [27]) gives
@@
This resembles a standard drift-diffusion equation if the second derivatives involving $t$ are neglected.
@@
Stochastic analyses originated by [17] and [9] (and further developed in subsequent works) demonstrate that relativity breaks Markovian processes in space, where one would expect that the set of spatial positions for the current particle state completely determine the next state [9, 10, 14]. In developing the relativistic theory, one might first consider bounding the particle speed in a Poisson-Kac process [20], which corresponds to the Telegrapher ($P_{1}$) equation and is non-Markovian [20, 10]. Spatially non-Markovian stochastic processes have been formulated in special [10, 30] and general [19] relativistic settings. Properly accounting for relativity in a collective sense, where a frame is moving relative to the frame in which diffusion is a good description, is a somewhat distinct problem. To this end, [14] examine the particle motion as a Poisson-Kac process along discrete directions, where the process exists in its usual form in one inertial (comoving) frame, finding in 1D that the effective diffusion coefficient in another inertial frame is $1/\gamma^{3}$ times the comoving coefficient; [14] observe that this follows the intuition that $D\sim\delta x^{2}/\delta t\sim(\delta x_{0}/\gamma)^{2}/(\gamma\delta t_{0})$, from Lorentz contraction of space and time dilation.
@@
For asymptotic analysis, the principal assumption is typically that the inverse of the macroscopic opacity is large compared to all length scales of interest [34]. Unlike the works above that focus on the statistical mechanics of the particles, this approach usually: starts from the full radiative transfer equation, imposes the assumed parameter scalings, and expands the radiation intensity in powers of a small-scale parameter (sometimes labeled $\varepsilon$). Asymptotic analysis has been applied to deriving a semi-relativistic diffusion equation [28, 26], where $\beta$ is assumed to be asymptotically small (equivalently: the hydrodynamic flow time scale is assumed much longer than the light crossing time at the length scale of interest). In the fully relativistic context, including general relativity, asymptotic analysis has been examined early on by [34] and subsequently generalized by [3] (see also [2]). Building off of these works, [31] demonstrate the truncated Thorne moment formalism (see [35]) converges to a diffusion equation with modifications from fluid acceleration and shear in the optically thick limit. The analyses of [3] and [31] consider both asymptotically large absorption and scattering opacity. This is consistent with the static-material scaling that furnishes equilibrium diffusion from non-linear thermal radiative transfer, but departs in not assuming that the light crossing distance on the time scale of interest is large compared to the spatial domain of interest [23].
@@
In the present study, we instead consider an asymptotic analysis that restricts to a special relativistic, 1D-planar, linear transport context, and we focus on the correspondence of the lab frame (some inertial frame) to the non-inertial frame comoving with fluid parcels of the background matter. In a departure from the work of [3] and [31] we attempt to use the standard parameter scaling for linear transport (see, for instance, [16, 22]) in the comoving frame. Thus the scattering opacity is assumed to be asymptotically large while the absorption opacity is assumed to be asymptotically small, and there is not an equilibrium solution (for instance, the Planck or Fermi-Dirac distributions in the comoving frame). In doing so, our objective is to: (i) highlight apparent non-physical pathologies that occur from applying this standard scaling when velocity gradients are present, (ii) demonstrate that a particular scaling furnishes the Poisson-Kac-derived $\gamma$-dressed result of [14], but for continuous directions and including the standard grey contribution from adiabatic expansion and Doppler shift (see, for instance [5]), (iii) compare solutions of the diffusion equation to a relativistic linear transport solution and better-known semi-relativistic diffusion equation [5].
@@
This paper is organized as follows. In Section 2, we present all considerations for deriving the fully relativistic 1D planar grey diffusion equation, starting from a linear transport equation with isotropic opacity and elastic isotropic scattering in the comoving frame. This includes considering the effect of frequency Doppler shift on integration over frequency for the lab-frame grey equation, in Section 2.1. We present properties of a simple class of functions, which are weighted powers of Lorentz transform (Doppler shift) factors of frequency, and their associated integrals in Section 2.2. These functions and their properties facilitate both the asymptotic derivation and the lab-frame harmonic expansion (Appendix AB). In Section 2.4, we present the development of the lab-frame asymptotic scaling that furnishes a closed comoving equation with respect to lab-frame space-time coordinates. This section is subdivided into sections that showcase the pathologies encountered from simpler asymptotic scaling attempts. In Section 3, we present a discretization of the fully relativistic diffusion equation, identifying terms with corresponding non- or semi-relativistic DDMC analogs. We then compare this discretization to lab-frame MC transport to provide a preliminary exploration of where the fully relativistic diffusion equation is applicable in Section 4. In the appendix, we provide: (A) details of the integrated Doppler function recursion, (B) the aforementioned harmonic ($P_{1}$) expansion, and finally (C) a brief non-relativistic attempt of the apparent opacity-only scaling of [34, 3, 31].
@@ 2.4 Asymptotic scaling
where $A[\cdot]$ is the original time-derivative operator (or the remainder of the transport equation), $B(\mu)$ is an as-yet unknown coefficient and $A_{\varepsilon}[\cdot]$ is a $\varepsilon$-scaled form of this operator. This scaling assumes that the residual excluded from the $\varepsilon$ coefficient scales as a leading-order contribution. The full set of scaling relations we use is
@@ 2.4.2 Scaling the comoving time derivative with the anisotropic term
We see that we have obtained a drift-diffusion equation for the leading-order scalar intensity using our lab-frame asymptotic scaling with $B(\mu)$ from Eq. (36). Furthermore, we note that we have not assumed a constraint on the value of $\beta$ itself. The diffusion stencil has three factors of $1/\gamma$, consistent with the finding of [14] for bi-directional 1D planar random walk. However, we have an additional $\beta$-dependent coefficient of $\phi_{0}^{(0)}$, which accounts for the energetic effect of Doppler shift and adiabatic expansion. To more clearly see this, we may neglect the time-dependence of $\beta$, and neglect terms of O($\beta^{2}$) and O($\beta\partial\beta/\partial x$), to obtain
@@
which is the standard semi-relativistic form (see, for instance, [5], Chapter 6). If $\beta$ is constant and approaches 1, assuming $q$ and $\phi_{0}^{(0)}$ remain bounded,
@@
which is the limiting behavior observed by [14].
@@ 2.4.3 Scaling the Lagrangian time derivative of $\beta$
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{Q_{0}}{\gamma}\;\;,$ (52)
where
$Q_{0}\equiv\frac{1}{2}\int_{-1}^{1}q_{0}(\mu_{0})d\mu_{0}\;\;.$
This equation now has the same coefficient for adiabatic and expansion effects as the standard semi-relativistic form, so retains mirrored behavior in scalar intensity when $(x,\beta)\rightarrow(-x,-\beta)$. In subsequent sections we examine Eq. (52), which corresponds to the following list of lab-frame scaling relations, now including Eq. (51),
@@ 3.1 Spatial stencil
Here we describe the discretization of Eq. (52) we use to explore properties of the solutions in Section 4. We solve the resulting discretized system deterministically, but we note that there exists a straightforward Discrete Diffusion Monte Carlo (DDMC) [13, 8, 1] interpretation of the discretized terms, and write them following the notation of [8]. In 1D, a grey deterministic solution is straightforward to implement, and we do not seek to hybridize lab-frame MC transport with DDMC in the scope of this work.
@@
$\eta_{i}\equiv\frac{4}{3}\left.\frac{\partial\beta}{\partial x}\right|_{i}$ (54)
@@
can be viewed as advection “leakage opacities” [8] (see [1] for an operator-split continuous treatment of advection).
The diffusion operator terms can be evaluated in the usual way in the domain interior [8],
@@
where $\pm$-superscripts again indicate evaluation just to the lower or upper side of the cell edge $i+1/2$. The formulation permits several options for discretizing the velocity field, for instance: if $\gamma$ is continuous then $\gamma_{i+1/2}^{-}=\gamma_{i+1/2}^{+}=\gamma_{i+1/2}$; if $\gamma$ is piecewise-constant, then one may set $\gamma_{i+1/2}^{-}=\gamma_{i}$ and $\gamma_{i+1/2}^{+}=\gamma_{i+1}$. Consequently, the diffusion leakage opacities, corresponding to the standard non-relativistic form [8], are
@@
Finally, to treat the boundary, we assume the asymptotic boundary layer condition used by [8] holds in the comoving frame,
@@
where $W(\cdot)$ is the angular weighting function for intensity incident just below $x_{i-1/2}$, $\psi_{0}^{-}$, and $\lambda$ is the extrapolation distance [16]. Equation (63) is consistent with applying scaling Eqs. (53) to the transport equation, keeping terms up to O($\varepsilon$), transforming intensity and opacity to the comoving frame, and then applying the boundary layer analysis of [16] (the factor of $\gamma$ from the transformation of $\sigma_{t}$ to $\sigma_{t,0}$ can be included in the transformation of the spatial coordinate to the stretched optical depth parameter). Equation (63) can also be obtained from the the static boundary condition specified in comoving space-time, assuming the comoving time derivative is 0. Following the discretization procedure of [8], the leakage opacity and boundary transmission probability are
@@ 3.2 Time stencil and stability
As the right panel of Fig. 1 suggests, increasing the time-centering parameter makes the time-evolution of the error more stable; for the parameters considered the stability profile of the drift-diffusion equation is comparable to that of the static diffusion equation. Furthermore, $\alpha=0.5$ should furnish 2nd-order accuracy in time. In Section 4, we consider time-independent code implementations for comparing accuracy of the diffusion equation to lab-frame Monte Carlo transport; thus we avoid the issue of the time-stability considered in the present section.
@@ 4 Numerical results
where $d_{p}\in i$ implies the sum is over the set of tracks of particle $p$ inside cell $i$. We note that expanding the quadratic and distributing the sum furnishes path length estimators for lab-frame energy density, flux and pressure in the context of the standard energy-momentum tensor transformation for comoving energy density (see, for instance, [5], Chapter 6).
For the following calculations, we fix the spatial domain $x\in[0,1]$ with 128 spatial cells. We also fix the comoving opacities as $\sigma_{t,0}=128$ and $\sigma_{a,0}=1/2$. All problems have vacuum boundary conditions on both sides of the spatial domain. Units are omitted as the linear transport and diffusion equations are spatially scale-free, so $\sigma_{t,0}=128$ and $\partial\beta/\partial x=0.03$ are 128 mean-free paths and a change in $\beta$ by 0.03 over some unit of length (e.g. the radius of a star). For the MC transport, we employ 512 particles per cell for the uniform volume source and 65,536 particles for the point source. All comparisons in this section use lab-frame time-independent code implementations of both the discrete diffusion and Monte Carlo.